40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 21 11:55:05 2015, Steve, Update, Optical Levers, 1103P noise measurement 1103P@qpdBS.jpgoplev1103P.pngoplevSum1103Pbs.png
    Reply  Tue Apr 21 15:03:27 2015, rana, Update, Optical Levers, 1103P noise measurement 
    Reply  Tue Apr 21 21:46:34 2015, rana, Update, Optical Levers, 1103P noise measurement oplev1103P.pngL1-ISI_SUS_ODC_FC496F_SPECTRUM-1113523216-86400.png
       Reply  Wed Apr 22 11:21:51 2015, Steve, Update, Optical Levers, BS & PRM oplevs are back to normal 
Message ID: 11233     Entry time: Wed Apr 22 11:21:51 2015     In reply to: 11232
Author: Steve 
Type: Update 
Category: Optical Levers 
Subject: BS & PRM oplevs are back to normal 

BS & PRM oplev is restored. Note: the F -150 lens was removed right after the first turning mirror from the laser. This helped Rana to get small spot on the qpd.

It also means that the oplev paths are somewhat different now.

 

 

 

ELOG V3.1.3-